# Shelf-Supervised Cross-Modal Pre-Training for 3D Object Detection

### v2

@@ Abstract
State-of-the-art 3D object detectors are often trained on massive labeled datasets. However, annotating 3D bounding boxes remains prohibitively expensive and time-consuming, particularly for LiDAR. Instead, recent works demonstrate that self-supervised pre-training with unlabeled data can improve detection accuracy with limited labels. Contemporary methods adapt best-practices for self-supervised learning from the image domain to point clouds (such as contrastive learning). However, publicly available 3D datasets are considerably smaller and less diverse than those used for image-based self-supervised learning, limiting their effectiveness. We do note, however, that such data is naturally collected in a multimodal fashion, often paired with images. Rather than pre-training with only self-supervised objectives, we argue that it is better to bootstrap point cloud representations using image-based foundation models trained on internet-scale image data. Specifically, we propose a shelf-supervised approach (e.g. supervised with off-the-shelf image foundation models) for generating zero-shot 3D bounding boxes from paired RGB and LiDAR data. Pre-training 3D detectors with such pseudo-labels yields significantly better semi-supervised detection accuracy than prior self-supervised pretext tasks. Importantly, we show that image-based shelf-supervision is helpful for training LiDAR-only and multi-modal (RGB + LiDAR) detectors. We demonstrate the effectiveness of our approach on nuScenes and WOD, significantly improving over prior work in limited data settings. Our code is available on GitHub.
@@ 1 Introduction
3D object detection is an integral component of the robot perception stack. To facilitate research in this space, the Autonomous Vehicle (AV) industry has released several large-scale 3D annotated multi-modal datasets [1, 2, 3]. However, LiDAR-based detectors pre-trained on nuScenes [1] cannot be easily applied to Argoverse [3] due to differences in sensor characteristics across hardware platforms. In practice, training 3D detectors for a new hardware platform requires re-annotating 3D bounding boxes at scale, which can be prohibitively expensive [4] and time consuming [5]. Instead, recent works [6, 7, 8, 9] demonstrate that self-supervised pre-training with unlabeled sensor data can improve downstream detection accuracy with limited labels. Notably, AVs already capture large-scale unlabeled multi-modal data localized to HD maps during normal testing [10], motivating the exploration of self-supervised pre-training for outdoor scenes.
@@
Bootstrapping 3D Representations with 2D Priors. VLMs [15, 16, 17] are trained on internet-scale data and show impressive zero-shot detection performance across many application domains [19]. Therefore, we posit that pre-training with noisy pseudo-labels distilled from vision-language models can outperform contrastive features learned from scratch (cf. Fig. 1). However, distilling image-based foundational knowledge to LiDAR remains an open challenge [20, 21, 22]. We address this challenge by explicitly “inflating” 2D instance masks to 3D bounding boxes via unlabelled multimodal (RGB + LiDAR) training data, using a careful combination of LiDAR cues, HD maps, and shape priors [10]. This allows us to produce accurate 3D pseudo-labels which can then be used to pre-train uni-modal or multi-modal 3D detectors. One significant advantage of our shelf-supervision over self-supervision is that it naturally produces a more aligned pretext task; instead of pre-training 3D detectors with contrastive learning, we learn from 3D pseudo-labels.
Contributions. We present three major contributions.
  1. 1.
    We propose Cross-Modal 3D Detection Distillation (CM3D), a zero-shot approach for generating 3D bounding boxes using off-the-shelf vision-language models.
  2. 2.
    We demonstrate that pre-training detectors with zero-shot 3D bounding box pseudo-labels achieves higher downstream detection accuracy than prior contrastive learning objectives.
  3. 3.
    We conduct extensive experiments to ablate our design choices and demonstrate that our simple approach achieves state-of-the-art semi-supervised 3D detection accuracy on the nuScenes benchmark.
@@ 2 Related Works
Unsupervised 3D Object Detection has gained significant interest in recent years as a way to auto-label large datasets without human annotators. Dewan et al. [23] introduced a model-free approach to detect and track the visible portions of objects by leveraging motion cues from LiDAR. More recently, Wong et al. [24] identified unknown instances through supervised segmentation and clustering [25]. However, both [23, 24] only reason about the visible extent of objects and are unable to generate amodal bounding boxes. Cen et. al. [26] used a supervised detector to generate bounding box proposals for unknown categories, and [27] exploited the correspondence between images and point clouds to generate object proposals. Similarly, Wilson et al. [10] employed an HD map and shape priors to “inflate” 2D detections to 3D cuboids. Utilizing priors such as surface normals [28, 29, 30] and motion cues [31, 32, 33, 34] has been shown to improve unsupervised object detection. You et al. [35] integrated cues from temporal scene changes (e.g. point ephemerality from multiple traversals) to detect mobile objects [36, 37]. Najibi et al. [38] used scene flow and clustering to generate 3D bounding boxes for moving objects. However, this method does not assign semantic labels to cuboids and cannot detect static objects. Similarly, [39, 40] use scene flow to cluster points and generate pseudo-labels. In contrast, CM3D uses 2D VLM detectors to localize, “lift”, and classify both static and moving objects.
@@
Self-Supervised Learning for 3D Representations has been applied to many domains, including object-centric point clouds [41, 42, 43], indoor scenes [44, 45, 46, 47, 48, 49], and outdoor environments [9, 7, 50]. Current techniques for self-supervised point cloud pre-training can be broadly characterized by their receptive field. Scene-based contrastive learning [6, 51] (often used in the context of representation learning for indoor scenes) may not capture the necessary fine-grained details required for recognizing small objects. In contrast, voxel-based contrastive learning (often used in the context of representation learning for outdoor scenes) inherently struggles with a restricted receptive field, limiting its ability to encode larger structures. Region-based representations [7, 52] attempt to strike a balance between coarse-grained scene representations and fine-grained voxel representations. However, most regions in outdoor scenes (e.g. ground plane and buildings) lack informative cues, making it difficult to learn robust features. More recently, multi-modal contrastive learning methods [8, 9] demonstrate that paired multi-modal sensors can provide complementary (self-)supervision for pre-training. PointContrast [6] established correspondences between different views of point clouds using a contrastive loss. DepthContrast [14] treated different depth maps as instances and discriminates between them. STRL [53] extracted invariant representations from temporally correlated frames in a 3D point cloud sequence. More recently, SimIPU [8] and CALICO [9] delve into multi-modal contrastive learning by utilizing paired RGB and LiDAR data. GeoMAE [50] frames point-cloud representation learning as a masked auto-encoding task, and uses geometric pretext tasks including occupancy, normals, and curvature estimation. In contrast to prior work, we argue that directly pre-training with noisy pseudo-labels from vision foundation models can be a more effective pre-training strategy for 3D object detection.
@@
Leveraging 2D Foundational Models for 3D Perception is an active area of research. Sautier et al. [54] introduced SLidR, a 2D-to-3D representation distillation method aimed at cross-modal self-supervised learning on large-scale point clouds. Mahmoud et. al. [55] subsequently extended SLidR with a semantically-aware contrastive constraint and a class-balancing loss. SEAL [56] takes inspiration from SLidR and used SAM [57] to generate 3D class-agnostic point cloud segments for contrastive learning. More recently, SA3D [58] extends SAM [57] to segment 3D objects with NeRFs. Anything-3D [59] combines BLIP [16], a pre-trained 2D text-to-image diffusion model, with SAM for single-view conditioned 3D reconstruction. 3D-Box-Segment-Anything utilizes SAM with a pre-trained 3D detector, VoxelNeXt [60], for interactive 3D detection and labeling. SAM3D [61] repurposes SAM to directly segment objects from BEV point clouds. Most recently, UP-VL [21] distilled 2D vision-language features [11] into LiDAR point clouds to generate amodal cuboids. We use foundational models to generate class-specific instance segmentation masks and “inflate” these masks to 3D cuboids using LiDAR cues, HD maps, and shape priors.
@@ 3 Cross-Modal 3D Detection Distillation (CM3D) with VLMs
Given a large unlabelled set of multi-modal (RGB + LiDAR) data, we combine 2D VLMs with 3D domain knowledge (given by 3D shape priors and map constraints) to generate 3D bounding box pseudo-labels for pre-training 3D detectors. We describe pseudo-label generation in Sec. 3.1 and our training pipeline in Sec. 3.2.
Figure 2: CM3D Pseudo-Label Pipeline. First, we prompt a vision-language detector (e.g. Detic [15]) with a class name (e.g. car) to generate 2D box proposals. Next, we prompt SAM [57] with the predicted 2D bounding boxes to generate high-quality instance segmentation masks. We then generate an oriented 3D cuboid using the set of LiDAR points that project to a given 2D instance mask. Specifically, we define the center of the cuboid to be the medoid of the LiDAR points, the dimensions (length, width, height) to be a fixed shape prior (similar to an anchor box) as reported by ChatGPT when prompted with the class name, and the orientation to be aligned with lane geometry provided from an HD map.
@@ 3.1 Generating 3D Bounding Box Pseudo-Labels
2D Mask Generation. The first stage of our pipeline requires producing accurate 2D instance masks for a fixed vocabulary of object categories. We prompt a VLM detector (e.g. Detic [15] or GroundingDINO [18]) with class names (e.g. car, bus, truck) to generate 2D box proposals. Though some VLM detectors (i.e., Detic) already produce instance segmentation masks, we found it more effective to prompt separate (foundational) models such as SAM [57] with the predicted 2D bounding boxes to generate high-quality segmentation masks. We evaluate the impact of using different VLMs in Appendix F.
@@
3D Projection. Next, we project the LiDAR points onto the 2D image plane and group all points within each instance mask. To produce a final 3D bounding box pseudo-label, we need to generate a 3D centroid, 3D orientation, and cuboid dimensions (width, length, height). Our key technical approach, described below and inspired by [10], is to leverage 3D priors in the form of maps (e.g. car cuboids should typically be aligned with map geometry) and shape priors (e.g. fine-grained categories such as trucks tend to have canonical dimensions).
@@
3D Cuboid Generation. We define an initial candidate 3D centroid to be the medoid of points within each mask, expressed as $m=\operatorname*{\arg\!\min}_{x\in P_{M}}\sum_{p\in P_{M}}{||p-x||_{2}}$, where $P_{M}$ is the set of LiDAR points that lie within the mask $M$. Note that this medoid tends to lie on the surface of the object visible to the LiDAR sensor rather than the true object centroid, which implies it will require refinement. To estimate a canonical width-height-length, we simply prompt an LLM (ChatGPT) with each class name and use the returned value for all instances. Although LLMs do not have access to our 3D training data, they have seen many descriptions of object shapes on the web and can provide reasonable prototypical 3D sizes of common objects (Appendix G). Lastly, we estimate box orientation for vehicles using lane geometry from an HD map and assign the direction of the nearest lane to each cuboid. We assign an orientation of zero degrees to all non-vehicles (e.g. pedestrians, barriers and traffic cones).
@@
3D Refinement. Fig. 2 visualizes our overall pseudo-label generation pipeline. Because many components of our pipeline provide only rough 3D estimates (e.g., not all trucks have precisely the same 3D dimension), we find that 3D labels can be refined via self-training, which is discussed in the next section. Appendix A ablates several strategies for improvement, including prompt engineering to improve VLM zero-shot 2D detections, LiDAR accumulation to improve medoid estimation, mask erosion to remove noisy LiDAR points near mask boundaries, medoid compensation to reduce object center estimation biases, and late-fusion of independent zero-shot 3D detectors to improve orientation and shape estimation. We find that imperfect pseudo-labels still provide an effective signal for pre-training.
@@ 3.2 Training with Pseudo-Labels
We use our generated pseudo-labels to pre-train 3D detectors that will be fine-tuned on a limited set of labeled data. We present a block diagram of our training pipeline in Fig 3. Notably, it is trivial to train any 3D detector with our pseudo-labels, making our proposed framework widely applicable. Further, although our pseudo-label generator requires paired RGB and LiDAR data for pseudo-label generation, we can easily train a LiDAR-only model like CenterPoint, which requires only LiDAR sweeps at inference.
In this paper, we evaluate CM3D by training two popular detectors, CenterPoint [62] and BEVFusion [63]. Following CALICO’s [9] experiment setup, we train BEVFusion with the CenterHead and turn off class-balanced grouping and sampling. Although copy-paste augmentation is widely used when training 3D detectors, we find that performance degrades with our noisy pseudo-labels (Appendix D). We train the CenterHead following standard 3D detection losses. We use the sigmoid focal loss (for recognition) [64] and L1 regression loss (for localization). Concretely, our loss function is as follows: $L=L_{HM}+\lambda L_{REG}$, where $L_{HM}=\sum_{i=0}^{C}SigmoidFocalLoss(X_{i},Y_{i})$ and $L_{REG}=|X_{BOX}-Y_{BOX}|$, where $X_{i}$ and $Y_{i}$ are the $i^{th}$ class’ predicted and ground-truth heat maps, while $X_{BOX}$ and $Y_{BOX}$ are the predicted and pseudo-label box attributes.
Self Training. Although directly pre-training with our pseudo-labels significantly improves mAP, it also yields lower NDS compared to prior work. We posit that this is due to the errors in size and orientation estimation, which can both be attributed to our simplistic 3D priors. Prior work suggests that self-training can be effective for semi-supervised learning [35, 21]; not only can it discover more objects [35, 21], but it may also refine noisy box labels. We find that self-training significantly improves orientation and shape estimates by exploiting the regularity of multimodal data; similar-looking objects will generate similar pseudo-labels due to the smoothness of the learned detection function. We ablate the impact of self-training in the Appendix B.
Figure 3: Pre-Training and Fine-Tuning Pipeline (a) First, we generate zero-shot 3D bounding boxes with CM3D as described in Fig. 2. (b) Next, we pre-train our 3D detector using the generated pseudo-labels. Notably, this process is no different from training 3D detectors on ground-truth labels. (c) Lastly, we fine-tune the detector using a small amount of labeled data.

### v3

@@ Abstract
State-of-the-art 3D object detectors are often trained on massive labeled datasets. However, annotating 3D bounding boxes remains prohibitively expensive and time-consuming, particularly for LiDAR. Instead, recent works demonstrate that self-supervised pre-training with unlabeled data can improve detection accuracy with limited labels. Contemporary methods adapt best-practices for self-supervised learning from the image domain to point clouds (such as contrastive learning). However, publicly available 3D datasets are considerably smaller and less diverse than those used for image-based self-supervised learning, limiting their effectiveness. We do note, however, that such 3D data is naturally collected in a multimodal fashion, often paired with images. Rather than pre-training with only self-supervised objectives, we argue that it is better to bootstrap point cloud representations using image-based foundation models trained on internet-scale data. Specifically, we propose a shelf-supervised approach (e.g. supervised with off-the-shelf image foundation models) for generating zero-shot 3D bounding boxes from paired RGB and LiDAR data. Pre-training 3D detectors with such pseudo-labels yields significantly better semi-supervised detection accuracy than prior self-supervised pretext tasks. Importantly, we show that image-based shelf-supervision is helpful for training LiDAR-only, RGB-only and multi-modal (RGB + LiDAR) detectors. We demonstrate the effectiveness of our approach on nuScenes and WOD, significantly improving over prior work in limited data settings. Our code is available on GitHub.
@@ 1 Introduction
3D object detection is an integral component of the robot perception stack. To facilitate research in this space, the Autonomous Vehicle (AV) industry has released several large-scale 3D annotated multi-modal datasets [1, 2, 3]. However, 3D detectors pre-trained on nuScenes [1] cannot be easily applied to Argoverse [3] due to differences in sensor characteristics across hardware platforms. In practice, training 3D detectors for a new hardware platform requires re-annotating 3D bounding boxes at scale, which can be prohibitively expensive [4] and time consuming [5]. Instead, recent works [6, 7, 8, 9] demonstrate that self-supervised pre-training with unlabeled sensor data can improve downstream detection accuracy with limited labels. Notably, AVs already capture large-scale unlabeled multi-modal data localized to HD maps during normal testing [10], motivating the exploration of self-supervised pre-training for outdoor scenes.
@@
Bootstrapping 3D Representations with 2D Priors. VLMs [15, 16, 17] are trained on internet-scale data and show impressive zero-shot detection performance across many application domains [18]. Therefore, we posit that pre-training with noisy pseudo-labels distilled from vision-language models can outperform contrastive features learned from scratch (cf. Fig. 1). However, distilling image-based foundational knowledge to LiDAR remains an open challenge [19, 20, 21]. We address this challenge by explicitly “inflating” 2D instance masks to 3D bounding boxes via unlabeled multimodal (RGB + LiDAR) training data, using a careful combination of LiDAR cues, HD maps, and shape priors [10]. This allows us to produce accurate 3D pseudo-labels which can then be used to pre-train uni-modal or multi-modal 3D detectors. One significant advantage of our shelf-supervision over self-supervision is that it naturally produces a more aligned pretext task; instead of pre-training 3D detectors with contrastive learning, we learn from 3D bounding box pseudo-labels.
Contributions. We present three major contributions. First, we propose Cross-Modal 3D Detection Distillation (CM3D), a zero-shot approach for generating 3D bounding boxes using off-the-shelf vision-language models. Notably, we show that pre-training detectors with 3D bounding box pseudo-labels achieves higher downstream detection accuracy than prior contrastive learning objectives. We conduct extensive experiments to ablate our design choices and demonstrate that our simple approach achieves state-of-the-art unsupervised and semi-supervised 3D detection accuracy on the nuScenes and WOD benchmarks.
@@ 2 Related Works
Unsupervised 3D Object Detection has gained significant interest in recent years as a way to auto-label large datasets without human annotators. Dewan et al. [22] introduced a model-free approach to detect and track the visible portions of objects by leveraging motion cues from LiDAR. More recently, Wong et al. [23] identified unknown instances through supervised segmentation and clustering [24]. However, both [22, 23] only reason about the visible extent of objects and are unable to generate amodal bounding boxes. Cen et al. [25] used a supervised detector to generate bounding box proposals for unknown categories, and Tian et al. [26] exploited the correspondence between images and point clouds to generate object proposals. Similarly, Wilson et al. [10] employed an HD map and shape priors to “inflate” 2D detections to 3D cuboids. Utilizing priors such as surface normals [27, 28, 29] and motion cues [30, 31, 32, 33] has been shown to improve unsupervised object detection. You et al. [34] integrated cues from temporal scene changes (e.g. point ephemerality from multiple traversals) to detect mobile objects [35, 36]. Najibi et al. [37] used scene flow and clustering to generate 3D bounding boxes for moving objects. However, this method does not assign semantic labels to cuboids and cannot detect static objects. Similarly, [38, 39] use scene flow to cluster points and generate pseudo-labels.
@@
Self-Supervised Learning for 3D Representations has been applied to many domains, including object-centric point clouds [40, 41, 42], indoor scenes [43, 44, 45, 46, 47, 48], and outdoor environments [9, 7, 49]. Current techniques for self-supervised point cloud pre-training can be broadly characterized by their receptive field. Scene-based contrastive learning [6, 50] (often used in the context of representation learning for indoor scenes) may not capture the necessary fine-grained details required for recognizing small objects. In contrast, voxel-based contrastive learning (often used in the context of representation learning for outdoor scenes) inherently struggles with a restricted receptive field, limiting its ability to encode larger structures. Region-based representations [7, 51] attempt to strike a balance between coarse-grained scene representations and fine-grained voxel representations. However, most regions in outdoor scenes (e.g. ground plane and buildings) lack informative cues, making it difficult to learn robust features. More recently, multi-modal contrastive learning methods [8, 9] demonstrated that paired multi-modal sensors can provide complementary (self-)supervision for pre-training. PointContrast [6] established correspondences between different views of point clouds using a contrastive loss. DepthContrast [14] treated different depth maps as contrastive instances and discriminates between them. STRL [52] extracted invariant representations from temporally correlated frames in a 3D point cloud sequence. More recently, SimIPU [8] and CALICO [9] delve into multi-modal contrastive learning by utilizing paired RGB and LiDAR data. GeoMAE [49] frames point-cloud representation learning as a masked auto-encoding task, and uses geometric pretext tasks including occupancy, normals, and curvature estimation.
@@
Leveraging 2D Foundational Models for 3D Perception is an active area of research. Sautier et al. [53] introduced SLidR, a 2D-to-3D representation distillation method aimed at cross-modal self-supervised learning on large-scale point clouds. Mahmoud et al. [54] subsequently extended SLidR with a semantically-aware contrastive constraint and a class-balancing loss. SEAL [55] takes inspiration from SLidR and used SAM [56] to generate 3D class-agnostic point cloud segments for contrastive learning. More recently, SA3D [57] extends SAM [56] to segment 3D objects with NeRFs. Anything-3D [58] combines BLIP [16], a pre-trained 2D text-to-image diffusion model, with SAM for single-view conditioned 3D reconstruction. 3D-Box-Segment-Anything utilizes SAM with a pre-trained VoxelNeXt [59] 3D detector for interactive labeling. SAM3D [60] repurposes SAM to directly segment objects from BEV point clouds. Most recently, UP-VL [20] distilled 2D vision-language features [11] into LiDAR point clouds to generate amodal cuboids.
@@ 3 Cross-Modal 3D Detection Distillation (CM3D) with VLMs
Given a large unlabelled set of multi-modal (RGB + LiDAR) data, we combine 2D VLMs with 3D domain knowledge (given by 3D shape priors and map constraints) to generate 3D bounding box pseudo-labels for pre-training 3D detectors (Fig. 2). Using these pseudo-labels, we can train LiDAR-only, RGB-only, or mulit-modal 3D detectors.
Figure 2: Overview. Given unlabeled, calibrated, and paired RGB images and LiDAR sweeps, we generate 3D bounding boxes psuedo-labels using foundational 2D priors from VLMs, map priors, and geometric clustering. We describe our unprojection step further in Fig. 3. Using these pseudo-labels, we can train LiDAR-only, RGB-only, or multi-modal 3D detectors.
@@ 3.1 Generating 3D Bounding Box Pseudo-Labels
2D Mask Generation. The first stage of our pipeline requires producing accurate 2D instance masks for a fixed vocabulary of object categories. We prompt a 2D VLM detector (e.g., Detic [15] or GroundingDINO [61]) with class names (e.g., car, bus, truck) to generate 2D box proposals. Although some VLM detectors (i.e., Detic) already produce instance segmentation masks, we found it more effective to prompt separate (foundation) models such as SAM [56] with the predicted 2D bounding boxes to generate high-quality segmentation masks. We evaluate the impact of using different VLMs in Appendix F. Notably, we find that Detic [15] produces better 3D pseudo-labels than GroundingDINO [61] in our CM3D pipeline.
@@
3D Projection. Next, we project the LiDAR points onto the 2D image plane and group all points within each instance mask. To produce a final 3D bounding box pseudo-label, we need to generate a 3D centroid, 3D orientation, and cuboid dimensions (width, length, height). Our key technical approach, described below and inspired by [10], is to leverage 3D priors in the form of maps (e.g., car cuboids should typically be aligned with map geometry) and shape priors (e.g., fine-grained categories such as trucks tend to have canonical dimensions).
@@
3D Cuboid Generation. We define an initial candidate 3D centroid to be the medoid of points within each mask, expressed as $m=\operatorname*{\arg\!\min}_{x\in P_{M}}\sum_{p\in P_{M}}{||p-x||_{2}}$, where $P_{M}$ is the set of LiDAR points that lie within the mask $M$. Note that this medoid tends to lie on the surface of the object visible to the LiDAR sensor rather than the true object centroid, which implies it will require refinement. To estimate a canonical width-length-height, we simply prompt an LLM (e.g., ChatGPT) with each class name and use the returned value for all instances. Although LLMs do not have access to our 3D training data, they have seen many descriptions of object shapes on the web and can provide reasonable prototypical 3D sizes of common objects (Appendix H). Lastly, we estimate box orientation for vehicles using lane geometry from an HD map and assign the direction of the nearest lane to each cuboid. We assign an arbitrary orientation to all non-vehicles (e.g., pedestrians, barriers and traffic cones).
@@
3D Refinement. Fig. 3 visualizes our overall pseudo-label generation pipeline. Because many components of our pipeline provide only rough 3D estimates (e.g., not all trucks have precisely the same 3D dimension), we find that 3D labels can be refined. Table 1 ablates several strategies for improvement, including prompt engineering to improve VLM zero-shot 2D detections, LiDAR accumulation to improve medoid estimation, mask erosion to remove noisy LiDAR points near mask boundaries, medoid compensation to reduce object center estimation biases, and late-fusion of independent zero-shot 3D detectors to improve orientation and shape estimation. We find that prompt engineering and medoid compensation provide the greatest improvement to pseudo-label quality. We present a detailed overview of each refinement step in Appendix B.
Figure 3: Unprojecting 2D Foundational Priors to 3D. First, we prompt an open-vocabulary 2D detector (e.g., Detic [15]) with a class name (e.g., car) to generate 2D box proposals. Next, we prompt SAM [56] with the predicted 2D bounding boxes to generate high-quality instance segmentation masks. We then generate an oriented 3D cuboid using the set of LiDAR points that project to a given 2D instance mask. Specifically, we define the center of the cuboid to be the medoid of the LiDAR points, the dimensions (length, width, height) to be a fixed shape prior (similar to an anchor box) as reported by ChatGPT when prompted with the class name, and the orientation to be aligned with lane geometry provided from an HD map.
@@ 3.2 Training with Pseudo-Labels
We use our generated pseudo-labels to pre-train 3D detectors that will be fine-tuned on a limited set of labeled data. We present a block diagram of our training pipeline in Fig 2. Notably, it is trivial to train any 3D detector with our pseudo-labels, making our proposed framework widely applicable. Further, although our pseudo-label generator requires paired RGB and LiDAR data for pseudo-label generation, we can easily train a LiDAR-only model like CenterPoint, which requires only LiDAR sweeps at inference. In this paper, we evaluate CM3D by training two popular detectors, CenterPoint [62] and BEVFusion [63]. Following CALICO’s [9] experiment setup, we train BEVFusion with the CenterHead and turn off class-balanced grouping and sampling. Although copy-paste augmentation is widely used when training 3D detectors, we find that performance degrades with our noisy pseudo-labels (Appendix E). We train the CenterHead following standard 3D detection losses. We use the sigmoid focal loss (for recognition) [64] and L1 regression loss (for localization). Concretely, our loss function is as follows: $L=L_{HM}+\lambda L_{REG}$, where $L_{HM}=\sum_{i=0}^{C}SigmoidFocalLoss(X_{i},Y_{i})$ and $L_{REG}=|X_{BOX}-Y_{BOX}|$, where $X_{i}$ and $Y_{i}$ are the $i^{th}$ class’ predicted and ground-truth heat maps, while $X_{BOX}$ and $Y_{BOX}$ are the predicted and pseudo-label box attributes.
Self Training. Although directly pre-training with our pseudo-labels significantly improves mAP, it also yields lower NDS compared to prior work. We posit that this is due to the errors in size and orientation estimation, which can both be attributed to our simplistic 3D priors. Prior work suggests that self-training can be effective for semi-supervised learning [34, 20]; not only can it discover more objects [34, 20], but it may also refine noisy box labels. We ablate the impact of self-training in the Appendix D. We find that even one round of self-training significantly improves average precision and reduces orientation estimation error.